OpenStudy (anonymous):

Simplify: (2+tan^2x/sec^2x)-1

OpenStudy (dumbcow):

rewrite the "1" as sec^2/sec^2 , then you can combine fractions Use identity: tan^2 = sec^2 - 1

OpenStudy (anonymous):

bur what if i change then into sine and cosine?

OpenStudy (anonymous):



OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yeah you could do that too it should simplify to

OpenStudy (dumbcow):

\[2 \cos^2 x + \sin^2 x - 1\]

OpenStudy (anonymous):

ok help me that way please :D

OpenStudy (dumbcow):

use identity: sin^2 + cos^2 = 1

OpenStudy (anonymous):